Circular No. 5948

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET MUELLER (1994c) Jean Mueller, Palomar Mountain Observatory, reports her discovery

of a comet in the course of the second Palomar Sky Survey. Confirmation has been made on films taken by K. Lawrence for E. Helin in the course of the Palomar Planet-Crossing Asteroid Survey.

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Mar. 10.46528 15^{h}51^{m}06.^{s}02 +2^{o}28′41′′.6 17  Mueller
10.50347 15 50 59.06 +2 29 08.4  _{”}
11.45538 15 47 57.71 +2 39 32.7 17.0 Lawrence
11.48090 15 47 52.20 +2 39 52.2  _{”}

J. Mueller (Palomar). 1.2-m Oschin Schmidt. Faint, diffuse nucleus with a

large, diffuse coma and no visible tail. K. Lawrence (Palomar). 0.46-m Schmidt. Diffuse without condensation.

1993 (243) 1 M. Belton, Solid State Imaging (SSI) Science Team Leader for the

Galileo project; and R. Carlson, Principal Investigator, Near Infrared Mapping Spectrometer (NIMS), report that a small object has been found in the close vicinity of (243) Ida. The object was detected in samples of SSI and NIMS data taken during Galileo’s flyby of Ida in Aug. 1993 and returned to the earth during 1994 Feb. 15–16. The detections were in data acquired on 1993 Aug. 29.693 and 29.697 UT (at the spacecraft), respectively. At the time, the spacecraft was 537.9 × 10^{6} km (3.596 AU) from the earth and 10 760 km from Ida. Seen in independent samples of NIMS and SSI data, the suspected satellite appears to have been ∼ 100 km from Ida and ∼ 10–20 times smaller. Further information, including approximate orbital characteristics, resolved images of the object, color and near infrared spectra, will be returned from the spacecraft over the next few months.

SUPERNOVA 1994C IN ANONYMOUS GALAXY A. Riess, P. Challis and R. Kirshner, Center for Astrophysics, report

that a spectrogram (range 350–910 nm) obtained at the Multiple Mirror Telescope confirms that this object is a supernova and suggests that it is of type Ia, 20 days past maximum. Measurements from absorption lines of the host galaxy give a redshift of 0.051.

1994 March 12 Brian G. Marsden